### v1

# Each of those eight coalition logics is also determined by six other kinds of models

Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions they meet. In this paper, we show that each of the eight coalition logics is also determined by the following six kinds of models, with the respective properties: single-coalition-first action models; single-coalition-first neighborhood models; clear grand-coalition-first action models; clear single-coalition-first neighborhood models; tree-like grand-coalition-first action models; tree-like single-coalition-first neighborhood models.
Keywords: coalition logics, action models, neighborhood models, seriality, independence of agents, determinism
1 Introduction
1.1 Eight coalition logics determined by eight classes of general concurrent game models
Coalition Logic $\mathsf{CL}$ proposed by Pauly ([10, 11]) is a central logic in logical research on strategic reasoning.
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{C}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{C}$ ensures $\phi$.
$\mathsf{CL}$ has two kinds of models.
The first kind is concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents who form coalitions; for every coalition, there is an availability function, which specifies available joint actions of the coalition at states; for every coalition, there is an outcome function, which specifies possible outcome states of joint actions of the coalition.
The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action in the model.
The second kind is alpha neighborhood models. Roughly, in an alpha neighborhood model: there are some states; there are some agents who form coalitions; for every coalition, there is an alpha neighborhood function, which specifies the alpha powers of the coalition. An alpha power is a scope into which the coalition can force the next moment to fall.
The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
[9] proposed that concurrent game models have three too-strong assumptions: the independence of agents, seriality, and determinism. Based on general concurrent game models, which do not have the three assumptions, a Minimal Coalition Logic $\mathsf{MCL}$ was presented.
Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some special kinds of situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. [8] showed the completeness of the eight logics in a uniform way. It should be pointed out that the completeness $\mathsf{CL}$ is already given in [11, 5] and the completeness of $\mathsf{MCL}$ is already given in [9].
General concurrent game models are called grand-coalition-first action models in this paper.
1.2 Our work
In this work, we show that each of the eight coalition logics is also determined by six other kinds of models with respective properties.
  • The first kind is single-coalition-first action models. In this kind of models, the outcome function for single coalitions determines the outcome functions for other coalitions and the availability functions for all coalitions.
    This kind of models makes good sense for situations where we know the space of future possibilities and how agents’ behavior changes the future beforehand.
  • The second kind is single-coalition-first neighborhood models. In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition. An actual power is a minimal scope in which the coalition can force the next moment to fall. In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
  • The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition do not have the same outcome.
  • The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the powers of an agent are pairwise disjoint.
  • The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history.
  • The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
The rest of the paper is structured as follows.
  • In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from the former to the latter.
  • In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models.
  • In Section 4, we discuss compositionality of availability functions and outcome functions in action models.
  • In Section 5: we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models, and show they can represent single-coalition-first action models.
  • In Section 6: we define clear grand-coalition-first action models, and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models, and show that they can represent clear grand-coalition-first action models.
  • In Section 7, we define tree-like grand-coalition-first action models, and show they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models, and show they can represent tree-like grand-coalition-first action models.
  • In Section 8, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties.
  • We make some concluding remarks in Section 9.
2 General settings of coalition logics
2.1 Language
Let $\mathtt{AP}$ be a countable set of atomic propositions and $\mathtt{AG}$ be a finite nonempty set of agents. The subsets of $\mathtt{AG}$ are called coalitions. $\mathtt{AG}$ is called the grand coalition.
Definition 1 (The language $\Phi$).
The language $\Phi$ is defined as follows, where $p$ ranges over $\mathtt{AP}$ and $\mathrm{C}\subseteq\mathtt{AG}$:
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$. The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual. Define the dual $[\langle\mathrm{C}\rangle]\phi$ of $\langle[\mathrm{C}]\rangle\phi$ as $\neg\langle[\mathrm{C}]\rangle\neg\phi$, indicating every available joint action of $\mathrm{C}$ enables $\phi$.
2.2 Action semantics
Let $\mathtt{AC}$ be a nonempty set of actions. For any $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$. For any $a\in\mathtt{AG}$, we write $\mathtt{JA}_{a}$ instead of $\mathtt{JA}_{\{a\}}$. We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
Definition 2 (Action models).
An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},$$\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an available action function for $\mathrm{C}$.
  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Different logics might put different constraints on action models.
Definition 3 (Action semantics of $\Phi$).
$\mathtt{AM},s\Vdash\top$ always
$\mathtt{AM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{AM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{AM},s\Vdash\phi$
$\mathtt{AM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
$\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
It can be verified:
$\mathtt{AM},s\Vdash[\langle\mathrm{C}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s,\mathrm{C})$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{AM},t\Vdash\phi$
2.3 Neighborhood semantics
Definition 4 (Neighborhood models and alpha neighborhood models).
A neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $\mathrm{C}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
A neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is called an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is closed under supersets.
Different logics may put different constraints on neighborhood models.
Definition 5 (Neighborhood semantics of $\Phi$).
$\mathtt{NM},s\Vdash\top$ always
$\mathtt{NM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{NM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{NM},s\Vdash\phi$
$\mathtt{NM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
$\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
The following fact indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing which formulas it satisfies.
Fact 1.
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\mathtt{AG},$$\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
Then, for every $s\in\mathtt{ST}$ and $\phi\in\Phi$, $\mathtt{NM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
2.4 Transformation of action semantics to neighborhood semantics
Definition 6 (Actual and alpha effectivity functions of action models).
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model, and $\mathrm{C}\subseteq\mathtt{AG}$.
The function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
The function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
$\mathtt{LE}_{\mathrm{C}}(s)=\{\mathrm{Y}\subseteq\mathtt{ST}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathrm{Y}\text{ for some }\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
In a sense, actual effectivity functions and alpha effectivity functions of action models contain important information on coalition powers.
Actual effectivity functions are more transparent in the following sense: each element of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions.
Definition 7 (Action models $z$-representable by neighborhood models and $\alpha$-representable by alpha neighborhood models).
We say a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $z$-represents an action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$.
We say a class of neighborhood models $\mathbf{NM}$ $z$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
We say an alpha neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $\alpha$-represents an action model $\mathtt{AM}$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{LE}_{\mathrm{C}}$.
We say a class of alpha neighborhood models $\mathbf{NM}$ $\alpha$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}$, and (2) every alpha neighborhood model in $\mathbf{NM}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
Intuitively, that a neighborhood model $\mathtt{NM}$ represents an action model $\mathtt{AM}$ means that $\mathtt{NM}$ contains all the information of $\mathtt{AM}$ to evaluate formulas of $\Phi$.
Theorem 1 (Representability implies equivalence).
  1. (1)
    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
  2. (2)
    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
Proof.
Here, we show only (1). We can show (2) in a similar way.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model such that $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$. Let $s\in\mathtt{ST}$ and $\phi\in\Phi$. We want to show $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
The proof is by induction of $\phi$.
It is trivial when $\phi$ is a proposition letter p, $\phi$ is $\top$, $\phi$ is $\neg\psi$ and $\phi$ is $\psi\wedge\theta$. We are done if we can show when $\phi$ is $\langle[\mathrm{C}]\rangle\psi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
$\Rightarrow$
Assume $\mathtt{AM},s\Vdash\phi$. Then, $\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$. Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$. Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$. Then, $\mathtt{NM},s\Vdash\phi$.
$\Leftarrow$
Assume $\mathtt{NM},s\Vdash\phi$. Then, $\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$. Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$. Then, $\mathrm{Y}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$. Then, $\mathtt{AM},s\Vdash\phi$.
3 Eight coalition logics determined by eight classes of grand-coalition-first action models
3.1 Eight classes of grand-coalition-first action models
Definition 8 (Grand-coalition-first action models).
A grand-coalition-first action model is a tuple $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathtt{AG}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 9 (Outcome functions and availability functions).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}.$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
Outcome functions in grand-coalition-first action models are given in a from-whole-to-part way: the outcome states of a coalition $\mathrm{C}$ performing a joint action $\sigma_{\mathrm{C}}$ at a state $s$ are determined by the outcome states of the grand coalition performing action profiles containing $\sigma_{\mathrm{C}}$ at $s$.
Fact 2.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\mid\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ for some }\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(s)\}$.
This fact indicates that availability functions in grand-coalition-first action models also are given in a from-whole-to-part way: available joint actions of a coalition $\mathrm{C}$ at a state are determined by available joint actions the grand coalition at the state.
Definition 10 (Successor functions).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Define the successor function $\mathtt{suc}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
For any set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
Fact 3.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathtt{suc}(s)$.
Proof.
$\Rightarrow$
Let $\mathrm{C}\subseteq\mathtt{AG}$. Let $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$. Then, there is $\sigma_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. That is $w\in\mathtt{suc}(s)$.
$\Leftarrow$
Let $w\in\mathtt{suc}(s)$. Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$.
Definition 11 (Seriality, independence, and determinism of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say:
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})}$ is a singleton.
We use $\mathrm{ES}$ to indicate the set of the eight strings: $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$. We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively. We use the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ to correspond to the eight combinations of the three properties, respectively.
For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
3.2 Eight axiomatic systems
Definition 12 (An axiomatic system for $\mathsf{MCL}$).
Axioms:
Tautologies ($\mathtt{A}\text{-}\mathtt{Tau}$): all propositional tautologies
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
Inference rules:
Modus ponens ($\mathtt{MP}$): $\dfrac{\phi,\phi\rightarrow\psi}{\psi}$
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$.
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$
where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
Definition 13 (Axiomatic systems for $\mathsf{MCL+X}$).
For all $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inferences of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [8]).
The axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas in $\Phi$ relative to the class of grand-coalition-first action $\mathtt{X}$-models.
Remarks
The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]). Are the other six classes of grand-coalition-first action models $\alpha$-representable? Are the eight classes of grand-coalition-first action models $z$-representable? It is not known yet.
4 About compositionality of availability functions and outcome functions in grand-coalition-first action models
We say availability functions of a grand-coalition-first action model is compositional if the availability function of a coalition can be derived from the availability functions of its members in a natural way. In a similar sense, we talk about compositionality of outcome functions of grand-coalition-first action models.
4.1 Neither outcome functions nor availability functions in grand-coa-lition-first action models are compositional
The following two facts are easy to verify, and we skip their proofs.
Fact 4.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
Then:
  1. (1)
    1. (a)
      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
    2. (b)
      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
  2. (2)
    1. (a)
      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
    2. (b)
      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, then $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
Note that none of the converses of these items holds. Figure 1 indicates a counter-example for the first two items and Figure 2 indicates a counter-example for the last two items.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{1}\atop\alpha_{2}\beta_{2}$$\alpha_{1}\beta_{2}\atop\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
Figure 1: An example of that outcome functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{2}$$\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
Figure 2: An example of that availability functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
4.2 Availability functions in grand-coalition-first action models meeting independence of agents are compositional
The following fact is easy to check:
Fact 5.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action $\mathtt{I}$-model.
Then:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$: $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$ if and only if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$: $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ if and only if $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
Remarks
Models of Coalition Logic are grand-coalition-first action $\mathtt{SID}$-models. In those models, availability functions are compositional, but outcome functions might not be. Figure 1 indicates a counter-example.
4.3 Outcome functions are compositional in many situations
There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. Outcome functions are compositional in these situations.
What follows are two examples, and the second is from [3].
Example 1.
Adam and Bob control a ship lock, which has two doors. Adam can close or open the front door and Bob can close or open the back door. They cannot open the two doors at the same time. This scenario can be represented by the grand-coalition-first action model depicted in Figure 3.
$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open_{b}})$$(\mathtt{skip},\mathtt{close_{b}})$$(\mathtt{open_{f}},\mathtt{skip})$$(\mathtt{close_{f}},\mathtt{skip})$
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1, where outcome functions are compositional. Here $c_{f}$ and $c_{b}$, respectively, express the propositions the front door is closed and the back door is closed. Here are some instances of outcome functions: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{2},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{2},\mathtt{skip})=\{s_{2}\}\cap\{s_{1},s_{2}\}=\{s_{2}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
Example 2.
A system has two processes, $a$ and $b$. The process $a$ can assign values $1$ or 00 to the variable $x$, and the process $a$ can assign values $1$ or 00 to the variable $y$. When $x=0$, then $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, then $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, then $b$ can leave the value of $y$ unchanged or change it to $1$. When $y=1$, then $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where outcome functions are compositional.
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$\{x_{1},y_{1}\}$$\{x_{1},y_{0}\}$$\{x_{0},y_{1}\}$$\{x_{0},y_{0}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{x:=1},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$
Figure 4: This figure indicates a grand-coalition-first action model where outcome functions are compositional. Here are some instances: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
The following settings about agency can be found in the literature, such as [6] and [7]: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.
5 Single-coalition-first action models and single-coalition-first neighborhood models
5.1 Single-coalition-first action models
Definition 14 (Single-coalition-first action models).
A single-coalition-first action model is a tuple $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathsf{out}_{a}:\mathtt{ST}\times\mathtt{JA}_{a}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $a$ such that for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 15 (Outcome functions and availability functions).
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model.
  • For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
  • For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
Clearly, the outcome functions in single-coalition-first action models are compositional.
Let $\mathtt{AC}$ be a set of actions and $\mathtt{JA}_{\mathrm{C}}$ be the set of joint actions of $\mathrm{C}$. For any $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}$, if for any $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma^{\prime}_{\mathrm{C}}|_{a}$.
Theorem 3.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model. The following three conditions are equivalent:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
    1. (a)
      for all $a\in\mathtt{AG}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$;
    2. (b)
      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$, where $\mathrm{C}\neq\emptyset$.
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$:
    1. (a)
      for all $a\in\mathtt{AG}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$;
    2. (b)
      $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
  3. (3)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
    1. (a)
      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
    2. (b)
      for all fusion $\sigma^{\prime\prime}_{\mathrm{C}}$ of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, for all $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$, then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. It suffices to show (2b).
Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{\mathrm{D}}$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. By (1b), $t\in\mathsf{out}_{b}(s,\sigma_{b})$, where $b\in\mathrm{D}$. Then, $t\in\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}$. Note $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{D}=\emptyset$. Similarly, we can show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}\neq\emptyset$ and $\mathrm{D}\neq\emptyset$.
Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\mathrm{D}=\{b_{1},\dots,b_{m}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$ and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$ and $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$. Then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$.
Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}\cup\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
(2) $\Rightarrow$ (1)
Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. It suffices to show (1b).
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$.
Note $\mathrm{C}=\{c_{1}\}\cup\mathrm{C}-\{c_{1}\}$. By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ = $\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$.
Note $\mathrm{C}-\{c_{1}\}=\{c_{2}\}\cup\mathrm{C}-\{c_{1},c_{2}\}$. By (2b), $\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$ = $\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2}\}})$.
Finally, we get
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}}(s,$ $\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}})$.
That is, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap$ $\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$.
Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
(2) $\Rightarrow$ (3)
Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
First, we show (3a).
  • Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. We want to show $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. We are done if we can show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. It suffices to show that there is a joint action $\sigma_{\overline{\mathrm{C}}}$ of $\overline{\mathrm{C}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
    Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\emptyset$. Then, $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there are $\sigma_{a_{1}},\dots,\sigma_{a_{n}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}}),\dots,w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
    Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, $w\in\bigcup\{\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\mid\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}\}$. Then, $w\in\mathtt{suc}(s)$. Assume ${\overline{\mathrm{C}}}=\emptyset$. Then, $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathtt{suc}(s)$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$. Assume ${\overline{\mathrm{C}}}\neq\emptyset$. Assume ${\overline{\mathrm{C}}}=\{b_{1},\dots,b_{n}\}$. As for every $b\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathtt{suc}(s)$, there are $\sigma_{b_{1}},\dots,\sigma_{b_{n}}$ such that $w\in\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}}),\dots,w\in\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{n}}$. By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
  • Let $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. We want to show $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. By (2b), $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$. Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
Second, we show (3b). Let $\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, and $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{\mathrm{C}}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, where $\sigma_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma_{c_{n}}\in\mathtt{JA}_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, where $\sigma^{\prime}_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma^{\prime}_{c_{n}}\in\mathtt{JA}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $1\leq i\leq n$.
Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{m}}(s,\sigma^{\prime}_{c_{m}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
(3) $\Rightarrow$ (2)
Assume (3). Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
First, we show (2a). Let $a\in\mathtt{AG}$. We want to show $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
  • Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$. By (3a), $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
  • Let $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. By (3a), $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$. Then, there is $\sigma_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. By (3a), $t\in\mathsf{out}_{a}(s,\sigma_{\mathtt{AG}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$.
Second, we show (2b). We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. By (3a), $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Note $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ and $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ but $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
Note $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We claim $\sigma_{\mathrm{C}}\nsubseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$. Assume $\sigma_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Note $\mathrm{C}\cap\mathrm{D}=\emptyset$. Then $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. There is a contradiction. Similarly, we can show $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(s)$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$ and $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}})$. Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ is a fusion of $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}$. By (3b), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. There is a contradiction.
Theorem 4.
Every single-coalition-first action model is a grand-coalition-first action model, but not vice versa.
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counter-example.
Remarks
Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1]. Convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
5.2 Single-coalition-first neighborhood models
[…]
Definition 16 (Single-coalition-first neighborhood models).
A single-coalition-first neighborhood model is a tuple $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST})-\{\emptyset\})$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a general cover of $\mathtt{suc}(s)$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 17 (Neighborhood functions).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the neighborhood function $\mathtt{nei}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathrm{C}=\emptyset\text{ and }\mathtt{suc}(s)\neq\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathrm{C}\neq\emptyset\end{cases}$
Fact 6.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
Proof.
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is clearly a general cover of $\mathtt{suc}(s)$.
Assume $\mathrm{C}=\emptyset$ and $\mathtt{suc}(s)\neq\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is clearly a general cover of $\mathtt{suc}(s)$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Let $t\in\mathtt{suc}(s)$. Then there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in X_{a_{1}}$, …, there is $X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in X_{a_{n}}$. Then $t\in X_{a_{1}}\cap\dots\cap X_{a_{n}}$. Note $X_{a_{1}}\cap\dots\cap X_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
Definition 18 (Seriality, independence, and determinism of single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say:
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if and only if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if and only if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if and only if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
As above, we let the eight strings in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
For any single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathtt{X}\in\mathrm{ES}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
Remarks
Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Models based on so-called tight alternating transition systems are single-coalition-first neighborhood $\mathtt{SID}$-models.
5.3 Representation of single-coalition-first action models by single-coalition-first neighborhood models
Theorem 5.
Every single-coalition-first action model is $z$-representable by a single-coalition-first neighborhood model.
This result is easy to show, and we skip its proof.
Theorem 6.
Every single-coalition-first neighborhood model $z$-represents a single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood $\mathtt{X}$-model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as follows:
  • $\mathtt{AC}=\{\alpha_{a-s-X}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{ and }X\in\mathtt{nei}_{a}(s)\}$.
  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-X}\in\mathtt{AC}$:
    $\mathsf{out}_{a}(s,\alpha_{x-u-X})=\begin{cases}X&\text{if $a=x$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
Let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
We claim that $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$.
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show that $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathrm{C}=\emptyset$.
Then, $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Note:
$\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{n}})$
$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}$
Let $X\in\mathtt{AE}(s,\mathrm{C})$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $X=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, where $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}},\dots,\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}(s,\sigma_{a_{n}})$ and $X$ is not empty. Then, $\sigma_{a_{1}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{a_{n}}\in\mathtt{av}_{a_{n}}(s)$.
Let $1\leq i\leq n$ and $\sigma_{a_{i}}=(a_{i},\alpha_{x-u-X_{i}})$. Note $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})\neq\emptyset$. Then, $\mathsf{out}(s,\sigma_{a_{i}})=X_{i}$, $\sigma_{a_{i}}=(a_{i},\alpha_{a_{i}-s-X_{i}})$ and $X_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $X_{1}\cap\dots\cap X_{n}=X$. Then, $X\in\mathtt{nei}(s,\mathrm{C})$.
Let $X\in\mathtt{nei}(s,\mathrm{C})$. Then, there are $X_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{n}\in\mathtt{nei}_{a_{n}}(s)$ such that $X_{1}\cap\dots\cap X_{n}=X$. Then, $\alpha_{a_{1}-s-X_{1}}\in\mathtt{AC},\dots,\alpha_{a_{n}-s-X_{n}}\in\mathtt{AC}$. Let $\sigma_{a_{1}}=\{(a_{1},\alpha_{a_{1}-s-X_{1}})\},\dots,\sigma_{a_{n}}=\{(a_{n},\alpha_{a_{n}-s-X_{n}})\}$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})=X_{1},\dots,\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=X_{n}$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=X_{1}\cap\dots\cap X_{n}=X$. Note $X$ is not empty. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$. Then, $X\in\mathtt{AE}_{\mathrm{C}}(s)$.
Theorem 7.
For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
First, it is easy to see the following statements are equivalent:
  • $\mathtt{S}\textbf{-}\mathtt{AM}$ is serial;
  • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
  • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}(s)\neq\emptyset$;
  • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$;
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial.
Second, it is easy to check the following statements are equivalent:
  • $\mathtt{S}\textbf{-}\mathtt{AM}$ is independent;
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{AE}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{AE}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent.
Third, it is easy to see the following statements are equivalent:
  • $\mathtt{S}\textbf{-}\mathtt{AM}$ is deterministic;
  • for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})$ is a singleton;
  • for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{AE}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton;
  • for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton;
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic.
Then, it follows that for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
Theorem 8 (Representation of single-coalition-first action models by single-coalition-first neighborhood models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of single-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of single-coalition-first neighborhood $\mathtt{X}$-models.
Proof.
Let $\mathtt{X}\in\mathrm{ES}$.
Let $\mathtt{S}\textbf{-}\mathtt{AM}$ be a single-coalition-first action $\mathtt{X}$-model. By Theorem 5, $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a single-coalition-first neighborhood $\mathtt{X}$-model.
Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a single-coalition-first neighborhood $\mathtt{X}$-model. By Theorem 6, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a single-coalition-first action $\mathtt{X}$-model.
6 Clear grand-coalition-first action models and clear single-coalition-first neighborhood models
6.1 Clear grand-coalition-first action models
Definition 19 (Clear grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Theorem 9.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, and let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
The following three conditions are equivalent:
  1. (1)
    for all $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
  3. (3)
    for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Assume (2) does not hold. Then, there is $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$.
Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
Then, there is $a\in\mathrm{C},\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$ and $\sigma_{a}\neq\sigma^{\prime}_{a}$. Then, $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
Then, $w\in\mathsf{out}_{a}(s,\sigma_{a})$ and $w\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. There is a contradiction.
(2) $\Rightarrow$ (3)
(3) is a special case of (2).
(3) $\Rightarrow$ (1)
Assume (3). Assume (1) does not hold. Then, there is $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
Let $w\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. There is a contradiction.
Theorem 10.
Every clear grand-coalition-first action model is a single-coalition-first action model, but not vice versa.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is essentially a single-coalition-first action model, we need to show:
  1. (1)
    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
  2. (2)
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
  3. (3)
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$,
    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
By Fact 5, (1) holds. Clearly, (3) holds. It remains to show (2).
Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
Assume $\mathrm{C}=\emptyset$. By the definition of $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathtt{suc}(s)$;
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
By Fact 4, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to find a single-coalition-first action model with a $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, $\sigma_{a}\neq\sigma^{\prime}_{a}$, and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. Then it is not a clear grand-coalition-first action model.
Remarks
Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models. As observed by Ågotnes and Alechina, the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
6.2 Clear single-coalition-first neighborhood models
Definition 20 (Clear single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$, and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
Theorem 11.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
The following two conditions are equivalent:
  1. (1)
    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
By Fact 6, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, let $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$.
Note $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$. Then $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Then, there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=X_{a_{1}}\cap\dots\cap X_{a_{n}}$, and there is $X^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=X^{\prime}_{a_{1}}\cap\dots\cap X^{\prime}_{a_{n}}$.
Then $w\notin X^{\prime}_{a_{i}}$ for some $1\leq i\leq n$. Note $w\in X_{a_{i}}$. Then $X_{a_{i}}\neq X^{\prime}_{a_{i}}$. Note $u\in X_{a_{i}}$ and $u\in X^{\prime}_{a_{i}}$. Then $X_{a_{i}}\cap X^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
(2) $\Rightarrow$ (1)
(1) is a special case of (2).
By 9, clear grand-coalition-first action models can be equivalently defined by three different ways, which respectively concern all agents, all coalitions and the grand coalition. It would be expected that the following condition is equivalent to the three conditions given in Theorem 11: for all $\mathtt{AG}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$. This is actually not the case. What follows is a counter-example.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, where:
  • $\mathtt{ST}=\{s_{0},s_{1},s_{2},s_{3}\}$;
  • $\mathtt{suc}(s_{0})=\{s_{1},s_{2},s_{3}\}$;
  • $\mathtt{nei}_{a}(s_{0})=\{\{s_{1},s_{2}\},\{s_{2},s_{3}\}\}$ and $\mathtt{nei}_{b}(s_{0})=\{\{s_{2}\},\{s_{1},s_{3}\}\}$.
Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\mathtt{suc}(s_{0})$. However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\mathtt{suc}(s_{0})$.
Remarks
Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that those models are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
6.3 Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models
Theorem 12.
Every clear grand-coalition-first action model is $z$-representable by a clear single-coalition-first neighborhood model.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model. ∎
Theorem 13.
Every clear grand-coalition-first neighborhood model $z$-represents a clear single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 8. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
From Theorem 12, Theorem 13, and Theorem 7, the following result follows:
Theorem 14 (Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of clear grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models.
7 Tree-like grand-coalition-first action models and tree-like single-coalition-first neigh-borhood models
7.1 Tree-like grand-coalition-first action models
Definition 21 (Histories in grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
Definition 22 (Tree-like grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history $\theta$ from $r$ to $s$.
Theorem 15.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
The following three conditions are equivalent:
  1. (1)
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$;
  2. (2)
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
  3. (3)
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
Proof.
(1) $\Rightarrow$ (3)
Assume (1). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $a\in\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is an $\mathtt{AG}$-history from $r$ to $r$. Assume $(r,\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{m},r)$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$. Then, $(r,\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{m}|_{a},r)$ is an $a$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$. Let $(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$ be the $a$-history from $r$ to $s$. For every $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{i}_{\mathtt{AG}})$. Then $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
    Let $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ be an $\mathtt{AG}$-history from $r$ to $s$. We claim $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},$ $s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are identical.
    Assume $n\neq m$. Let $a\in\mathtt{AG}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then they are different. We have a contradiction.
    Assume $\sigma^{i}_{\mathtt{AG}}\neq\lambda^{i}_{\mathtt{AG}}$ for some $1\leq i\leq n$. Then there is $a\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{a}\neq\lambda^{i}_{\mathtt{AG}}|_{a}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then, they are different. We have a contradiction.
    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$. Let $a\in\mathtt{AG}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then, they are different. We have a contradiction.
(3) $\Rightarrow$ (2)
Assume (3). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$. For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{m},s_{m})$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$. Let $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ be the $\mathtt{AG}$-history from $r$ to $s$. Then, $,(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},$$\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is an $\mathrm{C}$-history from $r$ to $s$.
    Assume $(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$ which is different from $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},$ $s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$.
    For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
    Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. We have a contradiction.
(2) $\Rightarrow$ (1)
It is easy to see that (1) is a special case of (2).
Theorem 16.
Every tree-like grand-coalition-first action model is a clear grand-coalition-first action model, but not vice versa.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model. We are done if we can show for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Let $s\in\mathtt{ST}$, $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, there are at least two paths $(s,\sigma_{\mathtt{AG}},w)$ and $(s,\sigma^{\prime}_{\mathtt{AG}},w)$ from $s$ to $w$. Then, there are at least two paths $(r,\dots,s,\sigma_{\mathtt{AG}},w)$ and $(r,\dots,s,\sigma^{\prime}_{\mathtt{AG}},w)$ from r to w. Note that $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like grand-coalition-first action model. Then, there is a contradiction.
It is easy to find a clear grand-coalition-first action model with a loop. Then, it is not tree-like.
Remarks
In the literature, Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models. They are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined here. It is required in their definition that for every state, there is a unique computation, which is a sequence of states, from the root to the state. Consequently, they are not clear.
7.2 Tree-like single-coalition-first neighborhood models
Definition 23 (Histories in single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
Definition 24 (Tree-like single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}$ $\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ from $r$ to $s$.
Theorem 17.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
The following two conditions are equivalent:
  1. (1)
    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$;
  2. (2)
    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $r$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$. For every $1\leq i\leq m$, let $Z_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $Z_{i}\subseteq\mathrm{Y}_{i}$ and $s_{1}\in Z_{i}$. Then, $(s_{0},Z_{1},s_{1},\dots,Z_{m},s_{m})$ is an $a$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$.
    Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$. It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$ is a unique $a$-history from $r$ to $s$.
    Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$.
    For every $1\leq i\leq n$, let $Z_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq Z_{i}$. Then, $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
    Let $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ be a $\mathrm{C}$-history from $r$ to $s$. We claim $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ and $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ are identical.
    For every $1\leq i\leq m$, let $E_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $D_{i}\subseteq E_{i}$. Then, $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$ is a $a$-history from $r$ to $s$.
    Assume $n\neq m$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
    Assume $Z_{i}\neq D_{i}$ for some $1\leq i\leq n$. Then, $Z_{i}\cap D_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,$ $\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
(2) $\Rightarrow$ (1)
It easy to see that (1) is a special case of (2).
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which are respectively concerned with all agents, all coalitions, and the grand coalition. It would be expected that the following condition is equivalent to the two conditions in Theorem 17: there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathtt{AG}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. The counterexample given after Theorem 11 also works here.
7.3 Representation of tree-like grand-coalition-first action models by tee-like single-coalition-first neighborhood models
Theorem 18.
Every tree-like grand-coalition-first action model is $z$-representable by a tree-like single-coalition-first neighborhood model.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorem 4, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. Then, it is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model. ∎
Theorem 19.
Every tree-like grand-coalition-first neighborhood model $z$-represents a tree-like single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 8. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
From Theorem 18, Theorem 19, and Theorem 7, the following result follows:
Theorem 20 (Representation of tree-like grand-coalition-first action models by tree-like single-coalition-first neighborhood models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of tree-like grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
8 Each of those eight coalition logics is determined by these six kinds of models, too
For every finite nonempty sequence $\chi$, we use $\mathsf{last}(\chi)$ to denote the last element of $\chi$.
Theorem 21.
Let $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model.
There is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ such that for all $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},$ $\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows:
  • $\mathtt{ST}^{\prime}=\mathtt{ST}_{0}^{\prime}\cup\mathtt{ST}_{1}^{\prime}\cup\mathtt{ST}_{2}^{\prime}\cup\dots$, where:
    • $\mathtt{ST}_{0}^{\prime}=\{w\}$;
    • $\mathtt{ST}_{k+1}^{\prime}=\{\chi-\sigma_{\mathtt{AG}}-u\mid\chi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
  • for every $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$:
    $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})=\{y\in\mathtt{ST}^{\prime}\mid y=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
  • for every $\chi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\chi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\mathsf{last}(\chi))\}$.
Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is what we want.
We show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is an $\mathtt{X}$-model.
[…]
  • Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent. Let $\chi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\chi)$. Then, $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\chi,\sigma_{\mathrm{D}})\neq\emptyset$.
    Then, there is $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$. Similarly, we can show $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\mathsf{last}(\chi))$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi))$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Let $u^{\prime}\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
    Then $u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma^{\prime}_{\mathtt{AG}})$ for some $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
    Then $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma^{\prime}_{\mathtt{AG}})$.
    Then, $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. Then, $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\chi)$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent.
  • Assume $(\mathtt{G}\textbf{-}\mathtt{AM},w)$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\chi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $s,s^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ where $s\neq s^{\prime}$. Let $s=\chi-\sigma_{\mathtt{AG}}-u$ and $s^{\prime}=\chi-\sigma_{\mathtt{AG}}-u^{\prime}$ where $u\neq u^{\prime}$. Then, $u,u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, there is a contradiction. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is deterministic.
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is a tree-like model.
We show for all $\chi$, and $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},w\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w\Vdash\phi$.
We put an induction on $\phi$. We consider only the cases $\phi=p$ and $\phi=\langle[\mathrm{C}]\rangle\psi$, and skip other cases.
  • $\phi$ is a proposition letter $p$.
    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash p$. Then, $p\in\mathsf{label}^{\prime}(\chi)$. Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$.
    $\Rightarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash p$. Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$. Then, $p\in\mathsf{label}^{\prime}(\chi)$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$.
  • $\phi$ is $\langle[\mathrm{C}]\rangle\psi$.
    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ such that for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
    We are done if we can show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$. Let $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$.
    Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\text{and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$.
    Then, $y\in\mathsf{out}^{\prime}(\chi,\sigma_{\mathrm{C}})$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$. By the inductive hyphothesis, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(y)\Vdash\psi$. That is $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    $\Rightarrow$
    Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
    Note $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$. Then $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$. Then $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$ is not empty. Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$.
    Let $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})\text{ and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then, there is $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ such that $t=\mathsf{last}(y)$.
    Then, $t\in\mathsf{out}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
Theorem 22.
For every $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
  1. (1)
    $\phi$ is valid with respect to the class of grand-coalition-first action $\mathtt{X}$-models;
  2. (2)
    $\phi$ is valid with respect to the class of single-coalition-first action $\mathtt{X}$-models;
  3. (3)
    $\phi$ is valid with respect to the class of single-coalition-first neighborhood $\mathtt{X}$-models;
  4. (4)
    $\phi$ is valid with respect to the class of clear grand-coalition-first action $\mathtt{X}$-models;
  5. (5)
    $\phi$ is valid with respect to the class of clear single-coalition-first neighborhood $\mathtt{X}$-models;
  6. (6)
    $\phi$ is valid with respect to the class of tree-like grand-coalition-first action $\mathtt{X}$-models;
  7. (7)
    $\phi$ is valid with respect to the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
Proof.
By Theorem 8 and Theorem 1, (2), and (3) are equivalent.
By Theorem 14 and Theorem 1, (4), and (5) are equivalent.
By Theorem 20 and Theorem 1, (6), and (7) are equivalent.
It suffices to show that (1), (2), (4) and (6) are equivalent. By Theorem 4, (1) implies (2). By Theorem 10, (2) implies (4). By Theorem 16, (4) implies (6). Suppose (6). Assume (1) does not hold. Then $\phi$ is false at a pointed grand-coalition-first action model $(\mathtt{G}\textbf{-}\mathtt{AM},w)$. Then, by Theorem 21, there is a tree-like grand-coalition-first model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ such that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)\nVdash\phi$. We have a contradiction. Then, (1).
9 Concluding remarks
Grand-coalition-first action models are those action models in which the outcome function for the grand coalition determines the outcome functions for other coalitions. Single-coalition-first action models are those action models in which the outcome functions for single coalitions determine the outcome functions for other coalitions. This class of models can be $z$-represented by the class of single-coalition-first neighborhood models. Clear grand-coalition-first action models are those single-coalition-first action models where different joint actions of a coalition have different outcomes. This class of models can be $z$-represented by the class of clear single-coalition-first neighborhood models. Tree-like single-coalition-first action models are those single-coalition-first action models where every state has a determined history. This class of models can be $z$-represented by the class of tree-like single-coalition-first neighborhood models. These kinds of models have different motivations.
Every kind of model has eight subclasses, determined by which of the properties seriality, independence of agents, and determinism hold. In this paper, we show that the subclasses of models that meet the same properties determine the same coalition logic.
Acknowledgments
Thanks go to Thomas Ågotnes, Valentin Goranko, and the audience of a logic seminar at Beijing Normal University.