### v2

# Each of those eight coalition logics is also determined by six other kinds of models

Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions are met. In this paper, we show that each of the eight logics is also determined by six other kinds of models, with respective properties, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
Keywords: coalition logics, action models, neighborhood models, seriality, independence of agents, determinism
1 Introduction
1.1 Eight coalition logics determined by eight classes of general concurrent game models
Coalition Logic $\mathsf{CL}$, proposed by Pauly ([10, 11]), is a central logic in logical research on strategic reasoning.
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{C}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{C}$ ensures $\phi$.
$\mathsf{CL}$ has two kinds of models.
The first kind is concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; for every coalition, there is an availability function, specifying available joint actions of the coalition at states; for every coalition, there is an outcome function, specifying possible outcome states of joint actions of the coalition. The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
The second kind is alpha neighborhood models. Alpha neighborhood models do not have actions. Consequently, they do not have availability functions and outcome functions. Instead, for every coalition, there is a neighborhood function that specifies the alpha powers of the coalition. An alpha power is a set of states into which the coalition can force the next moment to fall. The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
[9] argued that concurrent game models have three too strong assumptions: seriality, the independence of agents, and determinism. Based on general concurrent game models, which do not have the three assumptions, they proposed a Minimal Coalition Logic $\mathsf{MCL}$. Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some kinds of situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. [8] showed the completeness of the eight logics in a uniform way.
1.2 Our work
In this paper, we call general concurrent game models grand-coalition-first action models. In this work, we show that each of the eight coalition logics is also determined by six other kinds of models, with respective properties.
The first kind is single-coalition-first action models. In this kind of models, the outcome functions for single coalitions determine the outcome functions for other coalitions. This kind of models makes good sense for situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. The second kind is single-coalition-first neighborhood models. In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition. An actual power is a minimal scope into which the coalition can force the next moment to fall. In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition have different outcomes. Clear grand-coalition-first action models are quite common in game theory. The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the actual powers of an agent are pairwise disjoint.
The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history. The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
The rest of the paper is structured as follows. In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from action semantics to neighborhood semantics. In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models. In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models and show that they can represent single-coalition-first action models. In Section 5, we define clear grand-coalition-first action models and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models and show that they can represent clear grand-coalition-first action models. In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models and show that they can represent tree-like grand-coalition-first action models. In Section 7, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties. We conclude in Section 8.
2 General settings of coalition logics
2.1 Language
Let $\mathtt{AP}$ be a countable set of atomic propositions, and $\mathtt{AG}$ be a finite nonempty set of agents. The subsets of $\mathtt{AG}$ are called coalitions, and $\mathtt{AG}$ is called the grand coalition. In the sequel, when no confusion arises, we often write $a$ instead of $\{a\}$, where $a\in\mathtt{AG}$.
Definition 1 (The language $\Phi$).
The language $\Phi$ is defined as follows, where $p$ ranges over $\mathtt{AP}$ and $\mathrm{C}\subseteq\mathtt{AG}$:
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$. The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual.
2.2 Action semantics
Let $\mathtt{AC}$ be a nonempty set of actions. For every $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$. Joint actions of the grand coalition are called action profiles. In the sequel, we sometimes use sequences of actions to represent joint actions of coalitions, where an order among agents is implicitly presupposed. We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
For every $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{D}\subseteq\mathrm{C}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we use $\sigma_{\mathrm{C}}|_{\mathrm{D}}$ to indicate the restriction of $\sigma_{\mathrm{C}}$ to $\mathrm{D}$, which is in $\mathtt{JA}_{\mathrm{D}}$.
Definition 2 (Action models).
An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Different coalition logics might put different constraints on action models.
Definition 3 (Action semantics).
$\mathtt{AM},s\Vdash\top$
$\mathtt{AM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{AM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{AM},s\Vdash\phi$
$\mathtt{AM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
$\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
2.3 Neighborhood semantics
Definition 4 (Neighborhood models and alpha neighborhood models).
A neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $\mathrm{C}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
A neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
Different coalition logics might put different constraints on neighborhood models.
Definition 5 (Neighborhood semantics).
$\mathtt{NM},s\Vdash\top$
$\mathtt{NM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{NM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{NM},s\Vdash\phi$
$\mathtt{NM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
$\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
The following fact, which is easy to verify, indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing the set of formulas it satisfies.
Fact 1.
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
Then, for every $s\in\mathtt{ST}$ and $\phi\in\Phi$, $\mathtt{NM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
2.4 Transformation of action semantics to neighborhood semantics
Definition 6 (Actual effectivity functions and alpha effectivity functions of action models).
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
$\mathtt{LE}_{\mathrm{C}}(s)=\{\mathrm{Y}\subseteq\mathtt{ST}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathrm{Y}\text{ for some }\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
Actual effectivity functions are more transparent in the following sense: each element of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions.
Definition 7 (Action models $z$-representable by neighborhood models and $\alpha$-representable by alpha neighborhood models).
An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
A class of action models $\mathbf{AM}$ is $z$-representable by a class of neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
An action model $\mathtt{AM}$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
Intuitively, that a neighborhood model $\mathtt{NM}$ represents an action model $\mathtt{AM}$ means that $\mathtt{NM}$ contains all the information of $\mathtt{AM}$ to evaluate formulas of $\Phi$.
The following result can be easily shown:
Theorem 1 (Representability implies equivalence).
  1. (1)
    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
  2. (2)
    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}^{\alpha}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}^{\alpha}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.
3 Eight coalition logics determined by eight classes of grand-coalition-first action models
3.1 Eight classes of grand-coalition-first action models
Definition 8 (Grand-coalition-first action models).
A grand-coalition-first action model is a tuple $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathtt{AG}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 9 (Outcome functions and availability functions of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}.$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
Definition 10 (Successor functions of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},$ $\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Define the successor function $\mathtt{suc}$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$. It is easy to verify the following fact:
Fact 2.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathtt{suc}(s)$.
Fact 3.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
Note that neither of the converses of the two statements holds. Figure 1 indicates a counterexample for them.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts. A grand-coalition-first action model is indicated in the figure, where we suppose $\mathtt{AG}=\{a,b\}$. It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{a}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
Definition 11 (Seriality, independence, and determinism of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say:
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})}$ is a singleton.
We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively, and let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ correspond to the eight combinations of the three properties, respectively. We use $\mathrm{ES}$ to indicate the set of the eight strings.
For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
For every $\mathtt{X}\in\mathrm{ES}$, we use $\mathsf{MCL+X}$ to refer to the coalition logic determined by the class of grand-coalition-first action $\mathtt{X}$-models.
3.2 Eight axiomatic systems
Definition 12 (An axiomatic system for $\mathsf{MCL}$).
Axioms:
Tautologies ($\mathtt{A}\text{-}\mathtt{Tau}$): all propositional tautologies
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
Inference rules:
Modus ponens ($\mathtt{MP}$): $\dfrac{\phi,\phi\rightarrow\psi}{\psi}$
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$.
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
Definition 13 (Axiomatic systems for $\mathsf{MCL+X}$).
For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inference rules of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [8]).
For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas of $\mathsf{MCL+X}$.
Remarks
The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]). Are the other six classes of grand-coalition-first action models $\alpha$-representable? Are the eight classes of grand-coalition-first action models $z$-representable? All this is yet unknown.
4 Single-coalition-first action models and single-coalition-first neighborhood models
4.1 Outcome functions are compositional in many situations
We say outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
Outcome functions in grand-coalition-first action models are not compositional. Figure 2 gives a counterexample for this.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{1}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{2}$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
Figure 2: This figure offers a counterexample that outcome functions in grand-coalition-first action models are not generally compositional. We suppose $\mathtt{AG}=\{a,b\}$. Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted above, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$. In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$ In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$. Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$. This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. In these situations, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
What follows are two examples, and the second is from [3].
Example 1.
Adam and Bob control a ship lock with two doors: a front door and a back door. Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make the two doors open at the same time. This situation can be represented by the grand-coalition-first action model depicted in Figure 3, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open}\text{-}\mathtt{b})$$(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b})$$(\mathtt{open}\text{-}\mathtt{f},\mathtt{skip})$$(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip})$
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1. Here, $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed. It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action. For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{f})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
Example 2.
A system has two processes, $a$ and $b$. The process $a$ can assign $1$ or 00 to the variable $x$, and the process $b$ can assign $1$ or 00 to the variable $y$. When $x=0$, $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, $b$ can leave the value of $y$ unchanged or change it to $1$. When $y=1$, $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$\{x_{1},y_{1}\}$$\{x_{1},y_{0}\}$$\{x_{0},y_{1}\}$$\{x_{0},y_{0}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{x:=1},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$
Figure 4: This figure indicates a grand-coalition-first action model for the situation of Example 2. It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action. For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\mathtt{AG}}(s_{4},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{4},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{4},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
We want to mention that some literature, such as [6] and [7], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.
4.2 Single-coalition-first action models
Definition 14 (Single-coalition-first action models).
A single-coalition-first action model is a tuple $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathsf{out}_{a}:\mathtt{ST}\times\mathtt{JA}_{a}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $a$ such that for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 15 (Outcome functions and availability functions of single-coalition-first action models).
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$
Clearly, the outcome functions in single-coalition-first action models are compositional.
For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ equals to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
The following result offers two alternative definitions of sing-coalition-first action models.
Theorem 3.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
The following three sets of conditions are equivalent:
  1. (1)
    1. (a)
      for all $a\in\mathtt{AG}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$;
    2. (b)
      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
    It is easy to check this set of conditions is equivalent to the set of constraints on single-coalition-first action models given in Definition 14.
  2. (2)
    1. (a)
      for all $a\in\mathtt{AG}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$;
    2. (b)
      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
  3. (3)
    1. (a)
      for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
    2. (b)
      for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma_{\mathrm{C}}^{\prime},\sigma_{\mathrm{C}}^{\prime\prime}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). We only need to show (2b). Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\emptyset}(s,\emptyset)=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}\neq\emptyset$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. It suffices to show $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\emptyset}(s,\emptyset)$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $a\in\mathrm{D}$. By (1b), $t\in\mathsf{out}_{a}(s,\sigma_{\mathrm{D}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ $=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
The case that $\mathrm{C}\neq\emptyset$ and $\mathrm{D}=\emptyset$ can be handled similarly.
Assume $\mathrm{C}\neq\emptyset$ and $\mathrm{D}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, $\mathrm{D}=\{b_{1},\dots,b_{m}\}$, $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
(2) $\Rightarrow$ (1)
Assume (2). It suffices to show (1b). Let $\mathrm{C}$ be a nonempty subset of $\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
Note $\mathrm{C}=\{a_{1}\}\cup\mathrm{C}-\{a_{1}\}$. By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\mathsf{out}_{\mathrm{C}-\{a_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1}\}})$.
Note $\mathrm{C}-\{a_{1}\}=\{a_{2}\}\cup\mathrm{C}-\{a_{1},a_{2}\}$. By (2b), $\mathsf{out}_{\mathrm{C}-\{a_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1}\}})=\mathsf{out}_{a_{2}}(s,\sigma_{a_{2}})\cap\mathsf{out}_{\mathrm{C}-\{a_{1},a_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1},a_{2}\}})$.
Finally, we have $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
(2) $\Rightarrow$ (3)
Assume (2).
First, we show (3a). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. We want to show $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. It suffices to show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. By (2b), it suffices to show that there is $\sigma_{\overline{\mathrm{C}}}\in\mathtt{JA}_{\overline{\mathrm{C}}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
By (1a) and (1b), it is easy to verify $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Assume ${\overline{\mathrm{C}}}=\emptyset$. Clearly, what we want holds. Assume ${\overline{\mathrm{C}}}\neq\emptyset$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there is $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})$, …, there is $\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$ such that $w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
Let $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. By (2b), $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
Second, we show (3b). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma_{\mathrm{C}}^{\prime},\sigma_{\mathrm{C}}^{\prime\prime}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$. We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$. Let $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}=\emptyset$. It is easy to see $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $i$ such that $1\leq i\leq n$. By (1b), $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})$. By (1b), $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
(3) $\Rightarrow$ (2)
Assume (3).
First, we show (2a). Let $a\in\mathtt{AG}$. We want to show $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which equals to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
Let $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. By (3a), $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$. Then, there is $\sigma_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. By (3a), $t\in\mathsf{out}_{a}(s,\sigma_{\mathtt{AG}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$.
Second, we show (2b). Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. By (3a), $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Note $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ and $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Assume $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We want to get a contradiction.
We claim $\sigma_{\mathrm{C}}\nsubseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$. Assume $\sigma_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Note $\mathrm{C}\cap\mathrm{D}=\emptyset$. Then $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$, which is impossible. Similarly, we can show $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$ and $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}})$. Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ is a fusion of $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}$. By (3b), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We have a contradiction.
Theorem 4.
Every single-coalition-first action model is a grand-coalition-first action model, but not vice versa.
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counterexample.
Remarks
Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1]. It can be verified convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
4.3 Single-coalition-first neighborhood models
Let $\mathtt{ST}$ be a nonempty set of states and $\Delta_{1},\Delta_{2}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}$. Define $\Delta_{1}\odot\Delta_{2}=\{\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\mid\mathrm{Y}_{1}\in\Delta_{1},\mathrm{Y}_{2}\in\Delta_{2},\text{and }\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset\}$. Let $I$ be a nonempty set of indices. Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected. Note $\bigodot\{\Delta\}=\Delta$.
Definition 16 (Single-coalition-first neighborhood models).
A single-coalition-first neighborhood model is a tuple $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 17 (Neighborhood functions of single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the neighborhood function $\mathtt{nei}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}=\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}\neq\emptyset\end{cases}$
Fact 4.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
Proof.
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Let $t\in\mathtt{suc}(s)$. We want to show $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$ for every $a\in\mathtt{AG}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in\mathrm{X}_{a_{1}}$, …, there is $\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in\mathrm{X}_{a_{n}}$. Then $t\in\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$. Note $\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
Definition 18 (Seriality, independence, and determinism of single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say:
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
As above, we let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
For any $\mathtt{X}\in\mathrm{ES}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
Remarks
Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Goranko and Jamroga [4] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.
4.4 Representation of single-coalition-first action models by single-coalition-first neighborhood models
Theorem 5.
Every single-coalition-first action model is $z$-representable by a single-coalition-first neighborhood model.
This result is easy to show, and we skip its proof.
Theorem 6.
Every single-coalition-first neighborhood model $z$-represents a single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as follows:
  • $\mathtt{AC}=\{\alpha_{a-s-\mathrm{X}}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{and }\mathrm{X}\in\mathtt{nei}_{a}(s)\}$.
  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-\mathrm{X}}\in\mathtt{AC}$,
    $\mathsf{out}_{a}(s,\alpha_{x-u-\mathrm{X}})=\begin{cases}\mathrm{X}&\text{if $x=a$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
Let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
We claim that $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$.
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\}$.
Assume $\mathtt{suc}(s)=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset$ and $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
Let $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $\mathrm{X}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\neq\emptyset$, …, $\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\neq\emptyset$. Let $i$ be such that $1\leq i\leq n$ and $\sigma_{a_{i}}=\alpha_{x-u-\mathrm{X}_{i}}$. By the definition of $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})$, $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})=\mathrm{X}_{i}$, $x=a_{i}$, and $u=s$. Then $\sigma_{a_{i}}=\alpha_{a_{i}-s-\mathrm{X}_{i}}$ and $\mathrm{X}_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$. Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$ for some $\mathrm{X}_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{n}\in\mathtt{nei}_{a_{n}}(s)$. Then, $\alpha_{a_{1}-s-\mathrm{X}_{1}}\dots,\alpha_{a_{n}-s-\mathrm{X}_{n}}\in\mathtt{AC}$. Note $\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})=\mathrm{X}_{1}$, $\dots$, $\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{n}$. Let $\sigma_{\mathrm{C}}=\alpha_{a_{1}-s-\mathrm{X}_{1}}\cup\dots\cup\alpha_{a_{n}-s-\mathrm{X}_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}=\mathrm{X}$. Note $\mathrm{X}$ is not empty. Then, $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
Theorem 7.
For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for every $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$. Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then, for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
Let $\mathtt{X}\in\mathrm{ES}$. It is easy to see that the following three groups of statements are equivalent, respectively.
  1. (1)
    • $\mathtt{S}\textbf{-}\mathtt{AM}$ is serial;
    • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
    • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}(s)\neq\emptyset$;
    • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$;
    • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial.
  2. (2)
    • $\mathtt{S}\textbf{-}\mathtt{AM}$ is independent;
    • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
    • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{AE}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{AE}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
    • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
    • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent.
  3. (3)
    • $\mathtt{S}\textbf{-}\mathtt{AM}$ is deterministic;
    • for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})$ is a singleton;
    • for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{AE}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton;
    • for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton;
    • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic.
It follows that $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
The following result follows from Theorems 5, 6, and 7.
Theorem 8 (Representation of the class of single-coalition-first action $\mathtt{X}$-models by the class of single-coalition-first neighborhood $\mathtt{X}$-models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of single-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of single-coalition-first neighborhood $\mathtt{X}$-models.
5 Clear grand-coalition-first action models and clear single-coalition-first neighborhood models
5.1 Clear grand-coalition-first action models
Definition 19 (Clear grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
The following result offers two alternative definitions of clear grand-coalition-first action models.
Theorem 9.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
The following three conditions are equivalent:
  1. (1)
    for all $a\in\mathtt{AG}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
  3. (3)
    for all $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Suppose (2) does not hold. Then, there is $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$. Then, there is $a\in\mathrm{C}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$, and $\sigma_{a}\neq\sigma^{\prime}_{a}$. Let $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$, and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Then, $t\in\mathsf{out}_{a}(s,\sigma_{a})$ and $t\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$, which is impossible.
(2) $\Rightarrow$ (3)
(3) is a special case of (2).
(3) $\Rightarrow$ (1)
Assume (3). Suppose (1) does not hold. Then, there is $a\in\mathtt{AG}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. Let $t\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$, and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$, which is impossible.
For any set $\mathrm{X}$ of states, we say $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$ and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$. Note a general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
Fact 5.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
Proof.
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $t\in\mathtt{ST}$. Note $\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathtt{suc}(s)$. Let $\mathrm{X}_{1},\mathrm{X}_{2}\in\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ such that $\mathrm{X}_{1}\neq\mathrm{X}_{2}$. Then, $\mathrm{X}_{1}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{1})$ for some $\sigma_{\mathrm{C}}^{1}\in\mathtt{JA}_{\mathrm{C}}$ and $\mathrm{X}_{2}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{2})$ for some $\sigma_{\mathrm{C}}^{2}\in\mathtt{JA}_{\mathrm{C}}$. Note $\sigma^{1}_{\mathrm{C}}\neq\sigma^{2}_{\mathrm{C}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$, that is, $\mathrm{X}_{1}\cap\mathrm{X}_{2}=\emptyset$.
Note the other direction of the statement of this fact might not hold. Suppose $\mathtt{AG}=\{a\}$. Consider a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ such that:
  • $\mathtt{ST}=\{s\}$;
  • $\mathtt{AC}=\{\alpha_{1},\alpha_{2}\}$;
  • $\mathsf{out}_{a}(s,\alpha_{1})=\mathsf{out}_{a}(s,\alpha_{2})=\mathtt{ST}$.
It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$. However, $\mathtt{G}\textbf{-}\mathtt{AM}$ is not clear.
Theorem 10.
Every clear grand-coalition-first action model is a single-coalition-first action model, but not vice versa.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model, we need to show:
  1. (1)
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
  3. (3)
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$,
    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
By Fact 2, (1) holds. Clearly, (3) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
Assume $\mathrm{C}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{\mathtt{AG}}\}=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}=\mathtt{suc}(s)$;
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By Fact 3, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to find a single-coalition-first action model that is not clear.
Remarks
Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models.
As observed by Ågotnes and Alechina [12], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
5.2 Clear single-coalition-first neighborhood models
Definition 20 (Clear single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
The following result gives an alternative definition of clear single-coalition-first neighborhood models.
Theorem 11.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
The following two conditions are equivalent:
  1. (1)
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, we can assume there is $w$ such that $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$. It is easy to see $\mathtt{suc}(s)\neq\emptyset$. Then, $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$, which is a singleton. Note $\mathtt{nei}_{\mathrm{C}}(s)$ is not a singleton. Then, $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$, and there is $\mathrm{X}^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=\mathrm{X}^{\prime}_{a_{1}}\cap\dots\cap\mathrm{X}^{\prime}_{a_{n}}$. Note $w\notin\mathrm{Y}_{2}$. Then $w\notin\mathrm{X}^{\prime}_{a_{i}}$ for some $a_{i}$. Note $w\in\mathrm{X}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\neq\mathrm{X}^{\prime}_{a_{i}}$. Note $u\in\mathrm{X}_{a_{i}}$ and $u\in\mathrm{X}^{\prime}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\cap\mathrm{X}^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
(2) $\Rightarrow$ (1)
(1) is a special case of (2).
By Theorem 9, clear grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$. Actually, this is not the case. What follows is a counterexample.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
  • $\mathtt{ST}=\{s_{0},s_{1},s_{2},s_{3}\}$;
  • $\mathtt{suc}(s_{0})=\{s_{1},s_{2},s_{3}\}$;
  • $\mathtt{nei}_{a}(s_{0})=\{\{s_{1},s_{2}\},\{s_{2},s_{3}\}\}$ and $\mathtt{nei}_{b}(s_{0})=\{\{s_{2}\},\{s_{1},s_{3}\}\}$.
Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\mathtt{suc}(s_{0})$. However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\mathtt{suc}(s_{0})$.
Remarks
Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that they are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
5.3 Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models
Theorem 12.
Every clear grand-coalition-first action model is $z$-representable by a clear single-coalition-first neighborhood model.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model. ∎
Theorem 13.
Every clear single-coalition-first neighborhood model $z$-represents a clear grand-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
Theorem 14 (Representation of the class of clear grand-coalition-first action $\mathtt{X}$-models by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of clear grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models.
Proof.
Let $\mathtt{X}\in\mathrm{ES}$. Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action $\mathtt{X}$-model. By Theorem 12, $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood $\mathtt{X}$-model. Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a clear single-coalition-first neighborhood $\mathtt{X}$-model. By Theorem 13, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a clear grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a clear grand-coalition-first action $\mathtt{X}$-model.
6 Tree-like grand-coalition-first action models and tree-like single-coalition-first neighborhood models
6.1 Tree-like grand-coalition-first action models
Definition 21 (Histories in grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
Definition 22 (Tree-like grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
Note every tree-like model has a unique root.
The following result gives two alternative definitions of tree-like grand-coalition-first action models.
Theorem 15.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
The following three conditions are equivalent:
  1. (1)
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
  2. (2)
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
  3. (3)
    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
Proof.
(1) $\Rightarrow$ (3)
Assume (1). Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$. We want to show that there is an unique $\mathtt{AG}$-history from $r$ to $s$. Let $a\in\mathtt{AG}$.
Assume $s=r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$.
Assume there is an $\mathtt{AG}$-history $\theta_{\mathtt{AG}}$ from $r$ to $r$ which is different from $r$. It is impossible that $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. It is easy to check that $\theta_{a}=(s_{0},\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ is an $a$-history from $r$ to $r$. Clearly, $\theta_{a}$ is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $s\neq r$.
Let $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{a}=(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$.
For every $i$ such that $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\sigma^{i}_{\mathtt{AG}})$. Then, $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
Assume there is a different $\mathtt{AG}$-history $\theta^{\prime}_{\mathtt{AG}}$ from $r$ to $s$. Note it is impossible $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathtt{AG}}=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
We want to show $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for all $i$.
Assume $\sigma^{i}_{\mathtt{AG}}\neq\lambda^{i}_{\mathtt{AG}}$ for some $1\leq i\leq n$. Then, there is $b\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{b}\neq\lambda^{i}_{\mathtt{AG}}|_{b}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{b},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{b},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{b},t_{1},\dots,\lambda_{\mathtt{AG}}^{m}|_{b},t_{m})$ are $b$-histories from $r$ to $s$, which are different. We have a contradiction.
(3) $\Rightarrow$ (2)
Assume (3). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is an unique $\mathrm{C}$-history from $r$ to $s$.
Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$.
Assume there is a $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$ which is different from $r$. Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$.
For every $1\leq i\leq n$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then, $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$. We have a contradiction.
Assume $s\neq r$.
Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. Note it is impossible $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. It is easy to see $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Assume there is a different $\mathrm{C}$-history $\theta^{\prime}_{\mathrm{C}}$ from $r$ to $s$. Again, it is impossible $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathrm{C}}=(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$.
For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then, $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
It is easy to check that $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. We have a contradiction.
(2) $\Rightarrow$ (1)
(1) is a special case of (2).
Theorem 16.
Every tree-like grand-coalition-first action model is a clear grand-coalition-first action model, but not vice versa.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root. Let $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. It suffices to show $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. It is easy to check that both $(\theta_{\mathtt{AG}},\sigma_{\mathtt{AG}},t)$ and $(\theta_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}},t)$ are $\mathtt{AG}$-histories from $r$ to $t$, which are different. We have a contradiction.
It is easy to find a clear grand-coalition-first action model with a loop, which is not tree-like.
Remarks
Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section. For the former, it is only required that for every state, there is a unique computation, which is a sequence of states, from the root to the state.
6.2 Tree-like single-coalition-first neighborhood models
Definition 23 (Histories in single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
Definition 24 (Tree-like single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
The following result is easy to verify and we skip its proof.
Fact 6.
Every tree-like single-coalition-first neighborhood model is a clear single-coalition-first neighborhood model, but not vice versa.
The following result offers an alternative definition of tree-like single-coalition-first neighborhood models.
Theorem 17.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
The following two conditions are equivalent:
  1. (1)
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
  2. (2)
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume there is a different $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$. Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
Assume $\mathrm{C}=\emptyset$. Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\mathtt{suc}(s_{i-1})$. Let $a\in\mathtt{AG}$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$ for every $s\in\mathtt{ST}$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{m},s_{m})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $s\neq r$.
Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. It is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$. Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathtt{suc}(t_{0}),t_{1},\dots,\mathtt{suc}(t_{m-1}),t_{m})$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathtt{suc}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$. Then, $(t_{0},\mathrm{Z}_{1},t_{1},\dots,\mathrm{Z}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for all $0\leq i\leq n$. Then, $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$.
It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathrm{U}_{1},t_{1},\dots,\mathrm{U}_{m},t_{m})$. For every $1\leq i\leq m$, let $\mathrm{V}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{U}_{i}\subseteq\mathrm{V}_{i}$. Then, $(t_{0},\mathrm{V}_{1},t_{1},\dots,\mathrm{V}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for every $0\leq i\leq n$. Assume $\mathrm{Z}_{i}\neq\mathrm{U}_{i}$ for some $1\leq i\leq n$. Note $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood model. Then, $\mathrm{Z}_{i}\cap\mathrm{U}_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$, which is impossible. Then, $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
(2) $\Rightarrow$ (1)
(1) is a special case of (2).
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition is equivalent to the two conditions in Theorem 17: there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. The example given after Theorem 11 also works here.
6.3 Representation of tree-like grand-coalition-first action models by tee-like single-coalition-first neighborhood models
Theorem 18.
Every tree-like grand-coalition-first action model is $z$-representable by a tree-like single-coalition-first neighborhood model.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorems 16 and 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model. ∎
Theorem 19.
Every tree-like single-coalition-first neighborhood model $z$-represents a tree-like grand-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown there, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
The following result follows from Theorems 18, 19, and 7.
Theorem 20 (Representation of the class of tree-like grand-coalition-first action $\mathtt{X}$-models by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of tree-like grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
7 Each of those eight coalition logics is determined by these six kinds of models, too
Theorem 21.
For every $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
Proof.
Lert $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
In the following, for every finite nonempty sequence $\pi$, we use $\pi^{l}$ to denote the last element of $\pi$.
Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},$ $\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows:
  • $\mathtt{ST}^{\prime}=\mathtt{ST}_{0}^{\prime}\cup\mathtt{ST}_{1}^{\prime}\cup\mathtt{ST}_{2}^{\prime}\cup\dots$, where:
    • $\mathtt{ST}_{0}^{\prime}=\{s\}$;
    • $\mathtt{ST}_{k+1}^{\prime}=\{\pi-\sigma_{\mathtt{AG}}-u\mid\pi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},\text{and }u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
  • for every $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})=\{\pi-\sigma_{\mathtt{AG}}-u\mid u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
  • for every $\pi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\pi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\pi^{l})\}$.
Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$ and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is what we want.
We first make two claims:
  1. (1)
    for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$, if $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, then $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$;
  2. (2)
    for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, then there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
First, we show the claim (1). Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$. Assume $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. Note $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\subseteq\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $u\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, that is, $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
Second, we show the claim (2). Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $t\in\mathtt{ST}$. Assume $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$. Note $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\subseteq\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
We show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is an $\mathtt{X}$-model.
[…]
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent. Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$, and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\pi)$. To show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent, it suffices to show $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi)$. Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\pi,\sigma_{\mathrm{D}})\neq\emptyset$. By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}_{\mathrm{D}}(\pi^{l},\sigma_{\mathrm{D}})\neq\emptyset$. Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ and $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\pi^{l})$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\pi^{l})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\pi^{l},\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi)$.
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $\pi^{\prime},\pi^{\prime\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ such that $\pi^{\prime}\neq\pi^{\prime\prime}$. Let $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime}$ and $\pi^{\prime\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime\prime}$. Note $t^{\prime}\neq t^{\prime\prime}$. Then, $t^{\prime},t^{\prime\prime}\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. We have a contradiction.
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is tree-like.
We show for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$, which implies for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$. We put an induction on $\phi$. We consider only the case $\phi=\langle[\mathrm{C}]\rangle\psi$ and skip others.
Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ such that for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$. Let $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. By the claim (2), there is a $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ and for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. Note $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. Then $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$. Let $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. By the claim (1), $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{\prime l}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
Theorem 22.
For all $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
  1. (0)
    $\phi$ is valid with respect to the class of grand-coalition-first action $\mathtt{X}$-models;
  2. (1)
    $\phi$ is valid with respect to the class of single-coalition-first action $\mathtt{X}$-models;
  3. (2)
    $\phi$ is valid with respect to the class of single-coalition-first neighborhood $\mathtt{X}$-models;
  4. (3)
    $\phi$ is valid with respect to the class of clear grand-coalition-first action $\mathtt{X}$-models;
  5. (4)
    $\phi$ is valid with respect to the class of clear single-coalition-first neighborhood $\mathtt{X}$-models;
  6. (5)
    $\phi$ is valid with respect to the class of tree-like grand-coalition-first action $\mathtt{X}$-models;
  7. (6)
    $\phi$ is valid with respect to the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
Proof.
By Theorem 8 and Theorem 1, (1) and (2) are equivalent. By Theorem 14 and Theorem 1, (3) and (4) are equivalent. By Theorem 20 and Theorem 1, (5) and (6) are equivalent. It suffices to show that (0), (1), (3) and (5) are equivalent. By Theorem 4, (0) implies (1). By Theorem 10, (1) implies (3). By Theorem 16, (3) implies (5). By Theorem 21, (5) implies (1).
8 Conclusion
Li and Ju [9, 8] proposed eight coalition logics based on grand-coalition-first action models. In this work, we show that each of them is determined by six kinds of models, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
Acknowledgments
Thanks go to Thomas Ågotnes, Valentin Goranko, and the audience of a logic seminar at Beijing Normal University.